The length of a rectangle is three times the width, if the perimeter is 64 meters. Find the dimension of a rectangle. Show all work.
Answer by mananth(16949)

width = x
length = 3x
...
2*(l+w)=perimeter
2*(x+3x)=64
2*4x=64
8x=64
/8
x= 8 meters
length = 3x =3*8=24 meters
..
m.ananth@hotmail.ca